OpenStudy (anonymous):

when the base is 1why is the product always the same regardless of the exponent?

OpenStudy (mathmath333):

1^(anything) =1

OpenStudy (anonymous):

so i can put that? or do i need to make an equation

OpenStudy (mathmath333):

i asuumed it as log question \[\log_{1} 1=(any ~~number)\]

OpenStudy (mathmath333):

\[1^{28238282872828}=1\]

OpenStudy (anonymous):

thats the answer^

OpenStudy (mathmath333):

oh

OpenStudy (anonymous):

ohhh nvm its a word question idk its confusing

OpenStudy (anonymous):

im doing algebra online

OpenStudy (mathmath333):

cool

OpenStudy (anonymous):

yeah so the questions are kinda confusing

OpenStudy (anonymous):

i have another question

OpenStudy (anonymous):

its 6^2x8^2 can be written 6^2x8x2^3 why can the exponent 1 be removed?

OpenStudy (mathmath333):

u mean 8 ^1

OpenStudy (anonymous):

yeah i guess idk this is what it showed but yeah that sounds right

OpenStudy (mathmath333):

\(\large\tt \color{black}{ 6^2\times 8^2=6^2\times 8\times 8=6^2\times 8\times 2^3}\)

OpenStudy (anonymous):

yep thats what i have

OpenStudy (mathmath333):

u should learn exponents and its properties

OpenStudy (mathmath333):

http://www.mathsisfun.com/algebra/exponent-laws.html

OpenStudy (emmamink):

Because say we have 1^7. We can write that out and we get \[1\times1\times1\times1\times 1\times 1\times1\] It will always equal 1, no matter what power it is.

OpenStudy (anonymous):

im doing expanded form exponents

OpenStudy (anonymous):

and okay so i put that?

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

what about 6^2 x 8^2 can be written 6^2 x 8 x2^3. Why can the exponent 1 be removed?